A 20-year maturity bond with face value of $1,000 makes annual coupon payments and has a coupon rate of 6%.

a.

What is the bond’s yield to maturity if the bond is selling for $990? (Do not round intermediate calculations. Round your answer to 3 decimal places.)

  Yield to maturity %
b. What is the bond’s yield to maturity if the bond is selling for $1,000?
  Yield to maturity %
c.

What is the bond’s yield to maturity if the bond is selling for $1,190? (Do not round intermediate calculations. Round your answer to 3 decimal places.)

  Yield to maturity %
Answer #1

b. c. 6.088% la =RATE (20,6%*1000,-990, 1000) 6.000% b. =RATE(20,6%*1000,-1000, 1000) 4.535% c. =RATE(20,6%*1000,-1190, 1000)